Two charges, each 1 µC, are on the x axis, one at the origin and the other at x = 8 m. Find the electric field on the x axis at each of the following locations.
(a) x = -2 m
N/C

(e) At what point on the x axis is the electric field zero?
m
please tell me how you did this problem.
1
Expert's answer
2012-03-20T11:05:04-0400
a) E=E1+E2=q/(4*pi*eo*r1^2)+q/(4*pi*eo*r1^2)=q/(4*pi*eo) *
(1/r1^2+1/r2^2)=1*10^-9/(4*3.14*8.85*10^-12) * (1/4+1/(8+2)^2)=2.34
b)
E1-E2=0<=>
E1=E2<=>1/r1^2=1/r2^2<=>r1^2=r2^2<=>r1=r2<=>r=4m